Obtain the characteristic curves for the partial differential equation 4uxx+5uxy+uyy+ux+uy=0.
1
Expert's answer
2021-07-13T14:50:58-0400

Given Partial differential equation that is -


"=4u_{xx}+5u_{xy}+u_{yy}+u_{x}+u_{y}=0"


Now comparing this equation with the general form we get , which is -


"Au_{xx}+Bu_{xy}+Cu_{yy}+Du_{x}+Eu_{y}+Fu=G"



Now comparing both the equation we get ,


"A=4,B=5,C=1,D=1,E=1,F=0"



now taking out discriminant , which is equal to -


"B^{2}-4AC=25-4\times4\times1=9>0"


so the given characteristics curve will be hyperbolic .



Corresponding characteristics curve equation will be given as -


"\dfrac{dy}{dx}=\dfrac{-\xi_{x}}{\xi_{y}}" "=-(\dfrac{-B+\sqrt{B^{2}-4AC}}{2A})=-\dfrac{(-5+9)}{8}=-\dfrac{1}{2}"



"\dfrac{dy}{dx}=\dfrac{-\eta_{x}}{\eta_{y}}" "=-(\dfrac{-B-\sqrt{B^{2}-4AC}}{2A})=\dfrac{14}{8}=\dfrac{7}{4}"



To find "{\xi}" and "{\eta}" , we solve the above two equations for "y" , we get -


"y=\dfrac{-1}{2}x+c_{1}" and "y=\dfrac{7}{4}x+c_{2}"



Which give the value of two constants "c_{1}\ and \ c_{2}" as given below-


"c_{1}=y+\dfrac{1x}{2}" and "c_{2}=y-\dfrac{7x}{4}"


"{\therefore} {\xi}(x,y)=y+\dfrac{x}{2}=c_{1}" and "\eta({x,y})=y-\dfrac{7x}{4}=c_{2}"


These are equations of straight lines known as characteristics lines for given hyperbolic equation .